# Towards Interpreting Visual Information Processing in Vision-Language Models

### v1

@@ 1 Introduction
To understand the representations in the visual inputs for VLM and how the VLM processes them, we study LLaVA 1.5 7B (Liu et al., 2023a), a popular open-source VLM with competitive performance that uses CLIP (Radford et al., 2021) as an image encoder and Vicuna 13B (Chiang et al., 2023) as its language model. We also study LLaVA-Phi [^1] Sourced from the xtuner HuggingFace model. for our ablation experiments. We focus on a set of object identification tasks and our main contributions are as follows:
@@ 2.2.2 Processing of Object-Specific Representation
The processing of visual tokens may differ from text tokens. Soft prompts are very different from normal prompts: (Bailey et al., 2023), there is a modality gap between textual and visual representations in VLMs (Jiang et al., 2024). Furthermore, LLaVA is fine-tuned exclusively on visual question-answering (VQA) prompt-response pairs, without pre-training on next-token prediction. This may affect how the model processes visual information as compared to text. In Section 4.1, we attempt the logit lens method to see how the representations of visual tokens evolve through the layers.
@@ 3 Investigating the representations in mapped visual tokens
where $\bar{e}=\frac{1}{N}\sum_{i=1}^{N}e_{i}$ is the mean embedding across all visual tokens from 50,000 images from the ImageNet validation split (Deng et al., 2009). This replaces the hypothesized object-relevant tokens with an average token, effectively ablating their specific information. We then evaluate the impact of token ablation using three methods:
@@
    Generative Description: We prompt the model with “Describe the image” using both the original ($E_{A}$) and ablated embeddings ($E_{A}^{\prime}$). We then ascertain whether the target object $o$ is mentioned in the model’s generated description before and after ablation. If $o$ is mentioned with $E_{A}$ and not $E_{A}^{\prime}$, it indicates that the ablated tokens were crucial for the model to identify and include the object in its description.
@@
Results. Our results are presented in Table 1. We find that ablating the object tokens significantly impairs the model’s ability to identify the object. For comparable numbers of ablated tokens, object token ablation consistently results in larger performance decreases across all settings as compared to the gradient-based and random baselines. This suggests that the information about that object is localised to the region of the object token. Furthermore, hese findings are consistent across all three experimental settings and applicable to both LLaVA 1.5 and LLaVA-Phi.
Table 1: Performance degradation after token ablation. A lower percentage indicates a greater impact of the ablation, meaning the model is more likely to answer incorrectly, thus suggesting that the ablated tokens contain more localized object information. Average token counts apply to both the generative and polling settings for LLaVA 1.5. We find that ablating the object tokens with adjacent tokens (1 Buffer) significantly causes object identification performance to drop (in bold), more so than the integrated gradients and random baselines for a comparable number of tokens (in bold).
@@ 4.1 Analyzing the Residual Stream through Embedding Space
In some instances, okens correspond to their corresponding concepts in other languages, like in Figure 3(c) where the tokens correspond to “year” and “month” but in Chinese and Korean instead. This is somewhat surprising given that previous work has shown that Llama used for non-English languages tend to have English tokens in their intermediate activations when interpreted through the logit lens (Wendler et al., 2024).
@@
Figure 3: Examples of tokens and their positions that the logit lens yields in the late layers. In our labelling for an image, tokens that are part of a word are completed with an educated guess e.g. “diam”(ond)
@@
Case Study: Global Features. We sometimes find global-level features appearing in unexpected places, such numbers seemingly counting apples in background tokens (Figure 3(d)). When we tried ablating these specific visual tokens, the model’s description remained largely unchanged. Interestingly, the numbers persisted in some token positions in the logit lens even after ablation. These findings suggest that the global features revealed by the logit lens could be artifacts of the language model’s processing rather than direct information from the visual tokens. We hypothesise this to be one of the reasons as to why VLMs based on CLIP often have worse performance on image classification tasks than CLIP itself (Zhang et al., 2024).
@@ 4.2 Tracing Attention Flow
(3) For each window, we measure the decrease in accuracy and prediction probability for the correct token when blocking attention under three scenarios: (i) From object tokens and their surrounding buffer to the final token. (ii) From all non-object tokens to the final token. (iii) From all visual tokens except the last row to the last row of visual tokens. We conduct this experiment using LLaVA 1.5 on our curated Visual Question Answering (VQA) dataset of 100 images.
@@ 5 Related Work
Explainability. Classical techniques like GRAD-Cam (Selvaraju et al., 2017) and Integrated Gradients (Sundararajan et al., 2017), focus on explaining model outputs based on inputs. Recent tools like LVLM-Interpret (Stan et al., 2024) extend these methods to VLMs. While these approaches are valuable for understanding input-output relationships, they primarily focus on input attribution. Our work complements these by investigating the model’s internal processing mechanisms.
Mechanistic Interpretability (MI) for LLMs. Recent techniques such as causal tracing (Meng et al., 2022) and sparse coding methods (Bricken et al., 2023; Huben et al., 2023; Marks et al., 2024) have advanced our understanding of how LLMs perform tasks despite being primarily trained for next-token prediction. However, VLMs are not trained on next-token prediction for visual inputs (Liu et al., 2023b), raising questions about the generalizability of LLM processing mechanisms to visual inputs in VLMs. Our work extends some of these interpretability techniques to the multimodal domain to understand how they process visual inputs.
MI for Multimodal Models. In CLIP, (Goh et al., 2021) discovered multimodal neurons in the image encoder that respond to concepts across different modalities. Gandelsman et al. (2023; 2024) developed methods to decompose CLIP’s image representation, revealing property-specific roles for attention heads and polysemantic behavior in neurons. Rajaram et al. (2024) introduced a method to identify network subgraphs in CLIP that had interpretable functions. Rao et al. (2024); Daujotas (2024); Fry (2024) used sparse coding methods to yield interpretable features in the image encoder. However, CLIP’s architecture and training objectives differ substantially from generative VLMs, which our work focuses on.
For generative VLMs, Schwettmann et al. (2023) found neurons in the LM that corresponded to visual concepts, while Palit et al. (2023) adapted causal tracing for BLIP. While these studies provide valuable insights to specific components in the VLM, our work aims to investigate the representation in visual inputs and how those representations evolve in the LM.
@@ 6 Conclusion
In this work, we investigated the processing of visual information in Vision-Language Models, focusing on the LLaVA architecture. Our study reveals that object-specific information is highly localized within visual tokens corresponding to the object’s spatial location in the original image. This challenges the hypothesis that VLMs primarily rely on global image features encoded in register tokens.
Surprisingly, we found that representations at visual token positions evolve through the model’s layers to align with interpretable textual concepts, suggesting that VLMs may refine visual information towards language-like representations even without explicit pretraining on next-token prediction for visual inputs. Our attention blocking experiments further showed that VLMs directly extract object-specific information from relevant visual tokens in mid to late layers, while also processing broader contextual information from non-object tokens in earlier layers.
These findings advance our understanding of VLMs, bridging the gap between our knowledge of language and vision models. Our work provides a foundation for more interpretable and controllable multimodal systems. We hope to further investigate how our insights can be used to improve VLM performance, particularly in reducing hallucinations, model editing, and enhancing factual accuracy.
Limitations
We focused on LLaVA-type models, testing our findings on two variants (LLaVA 1.5 and LLaVA-Phi). We chose LLaVA for its simple architecture, viewing our findings as a starting point for interpreting more complex VLMs. However, our results may not generalize to VLMs with significantly different architectures or training procedures. Further research is needed to confirm whether similar mechanisms exist in other models.
Our study primarily investigates object identification tasks. We chose this focus as it provides a simpler, more direct setup compared to tasks requiring multi-step reasoning, such as multimodal factual association. While we attempted to diversify our experimental setup within this domain by testing the model’s ability to identify object through three distinct ways, our findings may not fully capture VLM behavior in more complex tasks like reasoning or open-ended question answering. The applicability of our results to these broader task types remains an open question.
@@
Clement Neo conceived the study, designed and performed the experiments, and led the writing of the manuscript. Luke Ong contributed to the writing and literature review. Philip Torr, Mor Geva and David Kreuger provided advice and comments. Fazl Barez served as the primary advisor for this work and helped significantly with the design and write-up of the paper.

### v2

@@ 1 Introduction
To understand the representations in the visual inputs for VLM and how the VLM processes them, we study LLaVA 1.5 7B (Liu et al., 2023a), a popular open-source VLM with competitive performance that uses CLIP (Radford et al., 2021) as an image encoder and Vicuna 13B (Chiang et al., 2023) as its language model. We also test on LLaVA-Phi [^1] Sourced from the xtuner HuggingFace model. and Qwen2-VL (Wang et al., 2024). We focus on a set of object identification tasks and our main contributions are as follows:
@@ 2.2.2 Processing of Object-Specific Representation
The processing of visual tokens may differ from text tokens. Soft prompts are very different from normal prompts (Bailey et al., 2023), and there is a modality gap between textual and visual representations in VLMs (Jiang et al., 2024). Furthermore, LLaVA is fine-tuned exclusively on visual question-answering (VQA) prompt-response pairs, without pre-training on next-token prediction. This may affect how the model processes visual information as compared to text. In Section 4.1, we attempt the logit lens method to see how the representations of visual tokens evolve through the layers.
@@ 3 Investigating the representations in mapped visual tokens
where $\bar{e}=\frac{1}{N}\sum_{i=1}^{N}e_{i}$ is the mean embedding across all visual tokens from 50,000 images from the ImageNet validation split (Deng et al., 2009). This replaces the hypothesized object-relevant tokens with an average token, effectively ablating their specific information. We do mean ablation to preserve the norm of the image token and keep them in-distribution, as their norms are typically much higher than the norm of text tokens (Bailey et al., 2023). We then evaluate the impact of token ablation using three methods:
@@
    Generative Description: We prompt the model with “Describe the image” using both the original ($E_{A}$) and ablated embeddings ($E_{A}^{\prime}$). We then ascertain whether the target object $o$ is produced verbatim in the model’s generated description before and after ablation. If $o$ is mentioned with $E_{A}$ and not $E_{A}^{\prime}$, it indicates that the ablated tokens were crucial for the model to identify and include the object in its description.
@@
Results. Our results are presented in Table 1. We find that ablating the object tokens significantly impairs the model’s ability to identify the object. For comparable numbers of ablated tokens, object token ablation consistently results in larger performance decreases across all settings as compared to the gradient-based and random baselines. This suggests that the information about that object is localised to the region of the object token. Furthermore, these findings are consistent across all three experimental settings and applicable to both LLaVA 1.5 and LLaVA-Phi.
Table 1: Performance degradation after token ablation. A lower percentage indicates a greater impact of the ablation, meaning the model is more likely to answer incorrectly, thus suggesting that the ablated tokens contain more localized object information. Average token counts apply to both the generative and polling settings for LLaVA 1.5. We find that ablating the object tokens with adjacent tokens (33.4 tokens on average, in bold) significantly causes object identification performance to drop, more so than the integrated gradients and random baselines for a comparable number of tokens (40 tokens, in bold).
@@ 4.1 Analyzing the Residual Stream through Embedding Space
To quantify this phenomenon, we analyzed 170 COCO validation images with objects of sizes between 20,000 and 30,000 square pixels (approximately 1/2 width and 1/3 height of the image). We found that in the best-performing layer for each image, an average of 23.7% of the object patch token positions correspond to the correct object class token. The best-performing layer occurs on average at layer 25.7 (out of 33 layers), confirming that mid-to-late layers tend to develop the strongest object-token correspondences (Figure 6).
@@
In some instances, tokens correspond to their corresponding concepts in other languages, like in Figure 3(c) where the tokens correspond to “year” and “month” but in Chinese and Korean instead. This is somewhat surprising given that previous work has shown that Llama used for non-English languages tend to have English tokens in their intermediate activations when interpreted through the logit lens (Wendler et al., 2024).
@@
Figure 3: Examples of tokens and their positions that the logit lens yields in the late layers. The labelled tokens come from a range of layers around the middle-to-late layers. In our labelling for an image, tokens that are part of a word are completed with an educated guess e.g. “diam”(ond)
@@
Testing Other Models. We extended our logit lens analysis to Qwen2VL-2B to verify if these findings generalize beyond the LLaVA family. Following the same methodology, we analyzed 253 COCO validation images and found that in the best-performing layer, an average of 6.5% of the visual tokens map to the correct object class token. The best-performing layer occurs at layer 25.1 (out of 29 layers), confirming the middle-to-late layer pattern observed in LLaVA (Figure 7). Interestingly, Qwen2VL shows non-zero correspondence even in early layers, possibly due to its cross-attention adapter providing better initial mapping compared to LLaVA’s simpler MLP adapter. Both models demonstrate increasingly aligned token representations with object tokens in later layers, suggesting this refinement pattern is a general property of adapter-style VLMs.
Case Study: Global Features. We sometimes find global-level features, such as numbers possibly referring to object counts, appearing in unexpected background tokens (Figure 3(d)). Ablating these specific tokens doesn’t change the model’s description, and the numbers persist in the logit lens even after ablation. This suggests these global features may be artifacts of the language model’s processing rather than direct information from visual tokens—potentially explaining why CLIP-based VLMs often underperform CLIP itself on classification tasks (Zhang et al., 2024).
@@ 4.2 Tracing Attention Flow
(3) For each window, we measure the decrease in accuracy and prediction probability for the correct token when blocking attention under three scenarios: (i) From object tokens and their surrounding buffer to the final token. (ii) From all non-object tokens to the final token. (iii) From all visual tokens except the last row to the last row of visual tokens [^2] We do this as recent work by Basu et al. (2024) suggests that the model may summarise visual information in the last row of visual tokens, though they study this in a different VQA task. If the model does use the last row of visual tokens as a ‘working summary’, then blocking the attention from all the visual tokens to the last row should prevent it from answering the question..
We conduct this experiment using LLaVA 1.5 on our curated Visual Question Answering (VQA) dataset of 100 images.
@@
Testing Other Models. To verify that our attention blocking results are not specific to LLaVA-1.5, we conducted additional experiments with LLaVA-Phi-3. Rather than using fixed layer groupings, we employed a sliding window of 10 layers to obtain finer-grained results across the network. We find that the results (Figure 8) closely mirror our original results: blocking attention from object tokens to the last token position shows the strongest effect in middle-to-late layers, while blocking attention from visual tokens to the last row continues to show minimal impact.
@@ 5 Related Work
Explainability. Classical techniques like GRAD-Cam (Selvaraju et al., 2017) and Integrated Gradients (Sundararajan et al., 2017), focus on input attribution, while recent tools like LVLM-Interpret (Stan et al., 2024) extend these methods to VLMs. Our work complements these by investigating the model’s internal processing mechanisms.
Mechanistic Interpretability (MI) for LLMs. Techniques such as causal tracing (Meng et al., 2022) and sparse coding (Bricken et al., 2023; Huben et al., 2023; Marks et al., 2024) have advanced our understanding of how LLMs perform tasks despite being primarily trained for next-token prediction, but VLMs are not trained on next-token prediction for visual inputs (Liu et al., 2023b). Our work extends some of these techniques to understand visual information processing in multimodal contexts.
MI for Multimodal Models. In CLIP, researchers have discovered multimodal neurons (Goh et al., 2021), decomposed image representations (Gandelsman et al., 2023; 2024), identified interpretable network subgraphs (Rajaram et al., 2024), and extracted interpretable features using sparse coding (Rao et al., 2024; Daujotas, 2024; Fry, 2024). In contrast, our work focuses on generative VLMs.
For generative VLMs, Schwettmann et al. (2023) found neurons corresponding to visual concepts, while Palit et al. (2023) adapted causal tracing for BLIP. While these studies help us understand the role of specific components in the VLM, we investigate the representations of the visual inputs.
@@
Parallel Work. Concurrent research by Jiang et al. (2025) independently arrived at similar logit lens findings and developed a hallucination reduction method based on these insights. Wu et al. (2024) observed comparable logit lens patterns across both visual and audio modalities.
@@ 6 Conclusion
Our investigation into visual information processing in LLaVA reveals that object information is highly localized within visual tokens corresponding to the object’s spatial location, challenging the hypothesis that VLMs primarily rely on global features in register tokens. Visual token representations evolve to align with interpretable textual concepts across layers, suggesting VLMs refine visual information toward language-like representations even without next-token prediction pretraining for visual inputs. Our attention blocking experiments suggest that VLMs may extract object information from relevant visual tokens in mid-to-late layers. These findings bridge the gap between our understanding of language and vision models, providing a foundation for more interpretable multimodal systems.
Limitations & Future Work. We focused on LLaVA-type models as a starting point for interpreting more complex VLMs, but our results may not generalize to significantly different architectures. Our study concentrates on object identification tasks, providing a simpler setup compared to multi-step reasoning. While we tested object identification in three ways, our findings may not fully represent VLM behavior in complex tasks like reasoning or open-ended question answering. Future work should explore practical applications in reducing hallucinations and enhancing factual accuracy, while examining whether these findings extend to broader task types and model architectures.
@@
Clement Neo conceived the study, designed and performed the experiments, and led the writing of the manuscript. Luke Ong contributed to the writing and literature review. Philip Torr, Mor Geva and David Krueger provided advice and comments. Fazl Barez served as the primary advisor for this work and helped significantly with the design and write-up of the paper.
@@ Appendix B Curated Images and Questions Examples
Appendix C Object Correspondences
Figure 6: Token-to-object correspondence in LLaVA-1.5 across transformer layers. Analysis of 170 COCO images shows peak correspondence (23.7%) occurs in later layers (average layer 25.7/33).
Figure 7: Token-to-object correspondence in Qwen2VL-2B showing increasing alignment in deeper layers, peaking at layer 25.1/29 with 6.5% correct mapping. Unlike LLaVA, Qwen exhibits non-zero correspondence in early layers.
Appendix D LLaVA-Phi-3 Blocking
Figure 8: Attention blocking results for LLaVA-Phi-3 using a sliding window of 10 layers. We find that similar to LLaVA-1.5, blocking attention from object tokens to the last token position has the strongest effect in middle-to-late layers, while blocking attention from visual tokens to the last visual row shows minimal impact throughout the network.